Show that fo = 1/2l[sqrt(T/u)] is homogenous
1
Expert's answer
2020-11-10T07:08:58-0500

Given equation is, "f_o = \frac{1}{2L}\sqrt{\frac{T}{\mu}}"

where "f_0" is frequency, L is the length, T is the tension and "\mu" is mass per unit length.


Equation is homogeneous only when dimensions on the both sides are eqaul.

Dimension of the left hand side, "[T^{-1}]"


Dimension of the right hand sides,

"\frac{1}{[L]}\sqrt{\frac{[MLT^{-2}]}{[ML^{-1}]}} = \frac{1}{[L]}\sqrt{[L^2T^{-2}]} = \frac{[LT^{-1}]}{[L]} = [T^{-1}]"



Since, dimension of the both sides are equal so the equation is homogeneous.